Briggs Company has income from operations of $76,356, invested assets of $303,000, and sales of $848,400. Use the DuPont formula to compute the return on investment. If required, round your answers to two decimal places.
(.25, .252, .3, .250 all were wrong)
Answer #1

Margin = Income from operation/Sales = 76356/848400 = 9%

Turnover = Sales/Average assets = 848400/303000 = 2.8

ROI = 9*2.8 = 25.2%